Average Vector Speed, Truck You stopped 8.4km while driving, moving 8.4km at a speed of 70km / h and stopped short of fuel. You left the vehicle and walked 30 minutes to the gas station 2.0 km away.
1
(a) Find the total transition from moving the truck to the gas station? (b) Find the interval Dr for the first time you move the truck to the gas station?
1
Expert's answer
2020-02-18T09:48:13-0500

a)


"s=8.4+2=10.4\ km"

b)


"T=\frac{8.4}{70}+0.5=0.62\ h=37.2\ min"

Average speed is


"v=\frac{10.4}{0.62}=16.8\frac{km}{h}"